OpenStudy (anonymous):

A vessel contains certain volume of milk and water mixed in the ratio 7 : 3. A part of this mixture is removed and replaced by an equal quantity of pure milk. Now the ratio of milk and water in the vessel is found to be 17 : 2. What proportion of the original volume has been replaced by pure milk?

OpenStudy (anonymous):

milk=7 water=3 changeinmilk=addedliquid-addedliquid*7/10 changeinwater=-addedliquid*3/10 (milk+changeinmilk)/(water+changeinwater)=17/2

OpenStudy (anonymous):

(7+.3x)/(3-.3x)=17/2

OpenStudy (anonymous):

8.222

OpenStudy (anonymous):

how added liquid in milk two time no removal of milk

OpenStudy (anonymous):

something went wrong here

OpenStudy (anonymous):

let T be the total volume, then (7/10)T is milk. Let x be the about that is replaced by pure milk. Then the new amount of milk is (7/10)(T-x) + x, which is 17/19 of the total. so, (7/10)(T-x) + x = (17/19)T x = (37/57)T about 64.91% of the total volume